OpenStudy (anonymous):

Please help Q4: Equation of Straight lines with slope and x-intercept equal...

OpenStudy (anonymous):

What exactly does this question mean to ask? If you could clarify any further..

OpenStudy (anonymous):

Equation of all straight lines with slope and x-intercept equal..?

ganeshie8 (ganeshie8):

y = m(x-c) slope is m, x intercept is c

OpenStudy (anonymous):

so , y=m(x-m)?

OpenStudy (anonymous):

Yes, that's pretty much it. Not a very interesting question, though, even for algebra..

ganeshie8 (ganeshie8):

you are given x-intercept of c, that means you have a point (c, 0 ) going through the line, and it has a slope of m. you can write the equation for a line with slope m, and going through a point (c, 0) y = 0 = m(x-c) y = m(x-c)